You are on a ski-lift 50 feet high. While the lift is stopped to let people off ,you accidentally knock your keys out of your pocket. They fall out of the seat towards the ground.
I'm suppose to write a vertical motion model for the path of the keys and find how long it'll take it to hit the ground.
I have 0= -16t^2 + vt + 50 but I don't know what to
Do next.

Answer by Alan3354(69443)

You are on a ski-lift 50 feet high. While the lift is stopped to let people off ,you accidentally knock your keys out of your pocket. They fall out of the seat towards the ground.
I'm suppose to write a vertical motion model for the path of the keys and find how long it'll take it to hit the ground.
I have 0= -16t^2 + vt + 50 but I don't know what to
Do next.
============================
h(t) = -16t^2 + vt + 50 is the model, or formula, for the height in feet vs time in seconds.
v is the initial vertical speed, = 0 in this case.
--> h(t) = -16t^2 + 50
Find the time when h(t) = 0
-16t^2 + 50 = 0
t^2 = 50/16
Use the positive value.
=====================
PS Ski lifts don't stop to let people off.